Menlo Company distributes a single product. The company’s sales and expenses for last month follow:

Total    Per Unit
  Sales $ 318,000 $ 20     
  Variable expenses 222,600 14     
  Contribution margin 95,400 $ 6     
  Fixed expenses 76,200
  Net operating income $   19,200

. What is the monthly break-even point in unit sales and in dollar sales?

Break even point in unit sales ? units
Break even point is sales dollars ?

2. Without resorting to computations, what is the total contribution margin at the break-even point?

Total contribution margin =

3-a. How many units would have to be sold each month to earn a target profit of $34,800? Use the formula method.

Units sold=

3-b. Verify your answer by preparing a contribution format income statement at the target sales level.

Menlo Company

Contribution Income Statement

total per unit
? ? ?
? ? ?
? 0 0
? ?
0

4. Refer to part 3 and now assume that the tax rate is 30%. How many units would need to be sold each month to an after-tax target profit of $34,800?  (Round the final answer to the nearest whole number.)

unit sales require = (in units)

5. Refer to the original data. Compute the company's margin of safety in both dollar and percentage terms. (Round your percentage answer to 2 decimal places (i.e .1234 should be entered as 12.34).)

Dollars Percentage
Margin of safety %

6. What is the company’s CM ratio? If monthly sales increase by $98,000 and there is no change in fixed expenses, by how much would you expect monthly net operating income to increase?

CM Ratio ? %
Net operating income increases by ?   
Answer #1

1) Break even unit = Fixed cost/Contribution margin per unit = 76200/6 = 12700 Units

Break even sales = 12700*20 = $254000

2) Total Contribution margin = $76200

3) Unit sold = (76200+34800)/6 = 18500 Units

3b) Contribution margin income statement

Total Per unit
Sales 370000 20
Variable cost 259000 14
Contribution margin 111000 6
Fixed cost 76200
Operating income 34800

4) Income before tax = 34800*100/70 = 49714

Units sold = (76200+49714)/6 = 20986 Units

5) Margin of safety

$ %
Margin of safety 318000-254000 = 64000 64000/318000 = 20.13%

6) CM ratio = 6/20 = 30%

Net operating income increase by 98000*30% = 29400